Solve : 

(D² + 5DD' + 5D'²) z = x sin (3x — 2y)


1
Expert's answer
2021-05-11T08:19:03-0400

We have given the differential equation,


"(D^2+5DD'+5D'^2)z = xsin(3x-2y)"


The auxiliary equation is


"m^2+5m+5 = 0"

"m = -5,0"

Hence, "CF = \phi_1(y-5x) + \phi_2(y)"

Particular integral PI can be calculated as :


"PI = \dfrac{1}{(D^2+5DD'+5D'^2)}x.sin(3x-2y)"


Factorize "\phi (D,D')" and apply these factors turn by turn


"PI = \dfrac{1}{D(D+5D')}xsin(3x-2y)"


"= \dfrac{1}{D} \int_{D+5D'}^{} xsin(3x-10x-2c_1)" because, "y-5x = c_1"


"= \dfrac{1}{D} \int xsin(-7x-2c_1)\delta x = \dfrac{1}{D} \int \dfrac{-xcos(-7x-2c_1)}{-7}+ \dfrac{sin(-7x-2c_1)}{49}"


"= \int_{D}^{}(\dfrac{x}{7}cos(3x-2y)+ \dfrac{1}{49} sin(3x-2y))\delta x"


"= \int_{D}^{}(\dfrac{x}{7}cos(3x-2c_2)+ \dfrac{1}{49} sin(3x-2y))\delta x" because "y= c_2"


"= \int_{D}^{}(\dfrac{x}{7}cos(3x-2c_2)\delta x+ \dfrac{1}{49} sin(3x-2y))\delta x"


Letting "a = \int_{D}^{}(\dfrac{x}{7}cos(3x-2c_2)\delta x"


"b = \dfrac{1}{49} sin(3x-2y))\delta x"


"a = \int_{D}^{}(\dfrac{x}{7}cos(3x-2c_2)\delta x = \dfrac{xsin(3x-2c_2)}{21} + \dfrac{1}{3}\dfrac{cos(3x-c_2)}{3}"


"a = \int_{D}^{}(\dfrac{x}{7}cos(3x-2c_2)\delta x = \dfrac{xsin(3x-2c_2)}{21} + \dfrac{1}{9}{cos(3x-c_2)}"


"b= \dfrac{1}{49} sin(3x-2y))\delta x"


"b = \dfrac{-cos(3x-c_2)}{49}+\dfrac{sin(3x-2c_2)}{147}"


Therefore,


"a+b = \dfrac{xsin(3x-2c_2)}{21} + \dfrac{1}{9}{cos(3x-c_2)}+ \dfrac{-cos(3x-c_2)}{49}+\dfrac{sin(3x-2c_2)}{147}"



The complete Integral "z=C.I.+P.I."


"z = \phi_1(y-5x) + \phi_2(y)+\dfrac{xsin(3x-2c_2)}{21} + \dfrac{1}{9}{cos(3x-c_2)}+ \dfrac{-cos(3x-c_2)}{49}+\dfrac{sin(3x-2c_2)}{147}"